# Propagation of Interplanetary Shocks in the Heliosphere [^1] July 17, 2024

### v1

@@ Abstract
Interplanetary shocks are one of the crucial dynamic phenomena in the Heliosphere. They accelerate particles to high energies, generate plasma waves, and can trigger geomagnetic storms in the terrestrial magnetosphere disturbing significantly our technological infrastructures. In this study, two IP shock events are selected to study the temporal variations of the shock parameters using magnetometer and ion plasma measurements of the STEREO--A and B, the Wind, Cluster fleet, and the ACE spacecraft. The shock normal vectors are determined using the minimum variance analysis (MVA) and the magnetic coplanarity methods. During the May 7, 2007 event, the shock parameters and the shock normal direction are consistent. The shock surface appears to be tilted almost the same degree as the Parker spiral, and the driver could be a Co–Rotating Interacting Region (CIR). During the April 23, 2007 event, the shock parameters do not change significantly except for the shock $\theta_{Bn}$ angle, however, the shape of the IP shock appears to be twisted along the perpendicular direction to the Sun-Earth line as well. The driver of this rippled shock is Stream Interaction Regions (SIRs)/CISs as well.
@@ 1 Introduction
The solar corona is hotter than the photosphere, the chromosphere, and the transient layers beneath it. As a result, the high temperatures ionize atoms, creating a plasma of free-moving electrons and ions, the so-called solar wind. Historically, (Parker, 1958) predicted the existence of the solar wind and coined terms describing it. He deduced it based on German astronomer Ludwig Bierman’s observation of how the comet tail always points away from the Sun (Biermann, 1957). The existence of the solar wind was confirmed by the Mariner 2 spacecraft (Snyder & Neugebauer, 1965). The solar wind is a collisionless plasma, whose speed is both supersonic and super-Alfvenic. The latter means that its speed exceeds the Alfvén speed, which is a representative speed of magnetohydrodynamic waves in a plasma. Due to abrupt change of the solar wind speed from supersonic to subsonic (e.g. by encountering an obstacle), a shock wave is emerged. Interplanetary (IP) shock waves are common in the heliosphere, which is a bubble-like region of space surrounding the Sun and extending far beyond the orbits of the planets and is filled with the solar wind. There are many varieties of shocks such as planetary bow shocks, shocks that are risen due to the stream interaction regions (SIR), which are called co-rotation interaction region (CIR) when extending beyond 1 AU (see for example in Koban et al., 2023), and coronal mass ejection (CME) driven shocks. IP shocks are one of the main and efficient accelerators of energetic particles in the heliosphere (Tsurutani & Lin, 1985; Keith & Heikkila, 2021). These accelerated particles can enter the geomagnetic field and are hazardous to astronauts and satellites. IP shocks driven by CMEs can trigger large geomagnetic storms that can damage oil and gas pipelines and interfere with electrical power infrastructures (Gonzalez et al., 1994). GPS navigation and high-frequency radio communications are also affected by ionosphere irregularities brought on by geomagnetic storms (Cid et al., 2014) and can cause internet disruptions around the world (Jyothi, 2021; Facsko et al., 2024). Therefore, IP shocks are important in determining and understanding space weather.
@@
Some dayside magnetospheric transient events (Zhang et al., 2022, see references therein) generated by the interaction of a discontinuity frozen in the solar wind and the terrestrial bow shock. This fact raises the questions of how stabile the discontinuities are, how long these phenomena travels with the solar wind, and whether they evolve during their travel, or show some wave dispersion (Facskó et al., 2005, István Ballai, personal communication). Dayside magnetospheric transient events are generated by tangential (TD) and rotational discontinuities (RD). However, there are catalogues for neither TDs nor RDs. Therefore, it is easier to study IP shocks because many were observed and listed previously (see Section 3).
@@
Reshetnyk et al. (2010) studied the 2D curvature of two IP shock events. Reshetnyk et al. (2010)’s paper showed that we do not know so much about the shape of discontinuities especially, the IP shocks. Facsko et al. (2010) extended the previuos study to 3D. More events and more discontinuities (IP shocks) were needed. A special automated shock identification method was developed (Facsko et al., 2025), however, finally, the appropriate events to study the temporal development and 3D shape of IP shocks were found in an open database developed by Nikbakhsh (2014); Kilpua et al. (2015). Finally, the origin of the IP shock are also very interesting. Based on out current knowledge the shock could be CME driven, the result of SIR/CIR development, or caused by a flare. The determination of the 3D shape of IP shocks with a backward projection might explain the origin of these very important events.
@@
The shape of an IP shock front depends on the inhomogeneities of the solar wind through the shock is propagating (Heinemann & Siscoe, 1974; Odstrčil & Pizzo, 1999; Dryer et al., 2004). Russell et al. (1983a, b) studied IP shock using two- and four-spacecraft measurements assuming that the IP shocks were planar locally. Terasawa et al. (2005) observed $\mathrm{15-20}^{\circ}$ shock surface ripple using Geotail (Nishida, 1994) and Advanced Composition Explorer (ACE; Stone et al., 1998) data. Szabo (2005) reported discrepancies of IP shock normal directions observing the same IP shock using ACE and Wind (Lin et al., 1995; Wilson et al., 2021), or Wind and IMP–8 (Team, 1974) spacecraft. Lepping et al. (2003) determined IP shock normals using Wind and IMP–8 observations at 1 AU (Astronomical Unit $\approx 1.5\times 10^{8}\,\mathrm{km}$). The average radius of the IP shock curvature was a few million km. Neugebauer & Giacalone (2005) collected and analyzed 26 IP shocks between the Earth’s bow shock and $L_{1}$ Lagrangian point. The average radius of the IP shock curvatures was a bit larger than Lepping et al. (2003)’s result, however, it also was a few million km. This result agreed with previous simulation results (Neugebauer & Giacalone, 2005).
@@
The Helios probe (Howard, 1975) observed IP shocks to evolve and change their behaviour at various heliospheric distances in the inner heliosphere (Lai et al., 2012) and between 1 and 5 AU by the Ulysses mission (Wenzel et al., 1992; Burton et al., 1992; Zhao et al., 2018). Trotta et al. (2023) studied shocklets using ACE, Wind, DSCOVR (Deep Space Climate Observatory), Time History of Events and Macroscale Interactions during Substorms (THEMIS, Angelopoulos, 2008) B and C measurements. Trotta et al. (2024) observed a CME driven shock at 0.07 AU and 0.7 AU using Parker Solar Probe (PSP, Fox et al., 2016), the Solar Orbiter (SolO, Marsch et al., 2005) observations. The PSP crossed a thick magnetosheath at the flank according to the early evolution phase of the CME. The SolO observed high energy (2 MeV) shock accelerated protons and a much more disturbed shock environment than the one observed by PSP. Therefore, the local small-scale environment of the shock was very different at different locations of the heliosphere.
In this study, we focus on the macroscopic structure and parameters of the observed IP shock events. The structure of this paper is as follows: we first list and describe briefly the missions, instruments, and databases that we used in this study in Section 2. We describe the applied analysis methods in Section 4. Then we introduce the datasets that we analysed in Section 5. We discuss and analyse the observations in Section 6. Finally, we summarize the results of our study in Section 7.
@@ 2.1 The STEREO mission
The PLAsma and SupraThermal Ion Composition (PLASTIC) instrument measures protons and alpha particle parameters, as well as the composition of heavy ions in the solar wind plasma (Galvin et al., 2008). The In-situ Measurements of Particles and CME Transients magnetic field experiment (IMPACT) suite of instruments consists of seven instruments among which three are located on on a 6-meter long boom while the other four are installed in the of the spacecraft (Acuña et al., 2008). The IMPACT measures the parameters of protons, heavy ions, and electrons, and the MAG magnetometer sensor in it measures the in situ magnetic fields in a range of ±$\pm 512\,nT$ with 0.1 nT accuracy (Kaiser & Adams, 2007).
We downloaded all data from NASA’s Coordinated Data Analysis Web (CDAWeb, https://cdaweb.gsfc.nasa.gov/). We obtained the STEREO--A and B magnetic field observation and the ion plasma data from the STEREO IMPACT with the time resolution of 100 ms and the STEREO PLASTIC instruments, respectively. The magnetic field and the plasma data of the STEREO--A and B are in the RTN spacecraft coordinate system, where R is radially outward from the Sun, T is along the planetary orbital, and N is northward direction.
@@ 2.2 The Wind mission
NASA’s Wind spacecraft was launched on November 1, 1994, (Lin et al., 1995; Wilson et al., 2021). The Wind was initially sent to $L_{1}$ Lagrange point, however, the reaching of $L_{1}$ was delayed in order to be able to study in advance the terrestrial the magnetosphere and lunar environment, as well. Following a sequence of orbital adjustments, the Wind spacecraft was positioned in a Lissajous orbit close to the $L_{1}$ Lagrange point in early 2004 for studying the incoming solar wind (NASA WIND team, 2020).
The spacecraft is equipped with eight instruments, however, we used only the Magnetic Field Investigation (MFI; Lepping et al., 1995) and the Three-Dimensional Plasma and Energetic Particle Investigation (3DP Lin et al., 1995) measurements in this study. The MFI consists of two magnetometers at the 12-meter boom measures magnetic field from 4 nT, to 65536 nT with a time resolution of 22 or 11 vectors per second in the calibrated high-resolution mode, while in the primary science mode the resolution can be three seconds, one minute and one hour (NASA WIND team, 2022). The 3DP measures the solar wind key parameters such as velocity, density, and temperature.
From CDAWeb, we got Wind magnetic and plasma data from the MFI and 3DP instrumens with a time resolution of 3 s, respectively. The magnetic field and the plasma data of the Wind are in the Geocentric Solar Ecliptic System (GSE) coordinate system, where X-axis is pointing to the Sun from the Earth, Y-axis is in the ecliptic plane showing opposite to the planetary motion, and Z-axis is the northward direction.
@@ 2.3 The ACE mission
NASA’s ACE spacecraft was launched on August 25, 1997 (Stone et al., 1998). The spacecraft is located at $L_{1}$ Lagrangian point similarly to the Wind spacecraft. The spacecraft is equipped with nine primary scientific instruments and one engineering instrument. We used in the study the measurements of the Magnetometer (MAG; Smith et al., 1998), and the Solar Wind Electron, Proton and Alpha Monitor (SWEPAM; McComas et al., 1998). The MAG consists of twin triaxial flux-gate magnetometers such that magnetometer sensors have between 3 and 6 vectors $s^{-1}$ resolutions for continuous observation of the interplanetary magnetic field (Smith et al., 1998).
@@ 2.4 The Cluster fleet
The FGM is composed of two tri-axial fluxgate magnetometers, which are installed on one of the two 5-meter radial booms. It measures in the dynamic range ±$\pm 65,536$ nT. At the highest dynamic level, the resolution is ±$\pm 8$ nT, and the time resolution is 100 vectors per second (Balogh et al., 1997, 2001; Organization, 2021). The CIS instrument measures three-dimensional ion distribution, and it is composed of two distinct sensors: the Composition Distribution Function (CODIF) sensor and the Hot Ion Analyzer (HIA) sensor (Rème et al., 1997, 2001). The CIS experiment is not operational for Cluster-2 and the HIA sensor is switched off for Cluster-4 due to a problem with the high voltage of the electrostatic analyzer (CIS team, 2021). Hence, for Cluster-2 and Cluster-4, the EFW instrument measurement became more indispensable for the analyses because we could obtain electron density in another way, independently from ion and electron plasma measurements (Gustafsson et al., 1997, 2001).
@@ 3 IP shock events and transformations
IP shock candidates are chosen from the shock lists in the Database of Heliospheric Shock Waves maintained at the University of Helsinki http://www.ipshocks.fi/. For the event selection, we chose the year 2007 because STEREO--A and STEREO--B were close to each other near the Sun-Earth line. In the first event, May 07 of 2007, the selected spacecraft are STEREO--A, STEREO--B, Wind, and the four Cluster satellites. For the second event, April 23, 2007, the spacecraft are STEREO--A and B, ACE, and Wind.
After choosing the shock candidates and obtaining the data, we transformeddata coordinates in such a way that all the coordinate systems are changed to the Heliocentric Earth Ecliptic (HEE) coordinate system, where the X-axis is toward the Earth from the Sun, Z-axis is the ecliptic northward direction, while Y axis completes the right-handed Cartesian system. This system is fixed with conserning the Sun-Earth line. To transform from the RTN and the GSE coordinate systems to the HEE coordinate system, we used Transformation de REpères en Physique Spatiale (TREPS; http://treps.irap.omp.eu/) online tool. The TREPS tool, which is developed by the French Plasma Physics Data Centre (CDPP), the national data centre in France for the solar system plasmas, is based on SPICE (Spacecraft, Planet, Instrument, C-matrix, and Events) information system kernels created by National Aeronautics and Space Administration (NASA)/Navigation and Ancillary Information Facility (NAIF) tool (Génot et al., 2018).

### v2

# Propagation of Interplanetary Shocks in the Inner Heliosphere [^1] December 4, 2024

@@ Abstract
Interplanetary shocks are one of the crucial dynamic phenomena in the Heliosphere. They can accelerate particles to high energies, generate plasma waves, and can trigger geomagnetic storms in Earth’s magnetosphere, significantly impacting technological infrastructure. In this study, two IP shock events are selected to study the temporal variations of the shock parameters using magnetometer and ion plasma measurements of the STEREO--A and B, the Wind, Cluster fleet, and the ACE spacecraft. The shock normal vectors are determined using the minimum variance analysis (MVA) and the magnetic coplanarity methods. During the May 7, 2007, event, the shock parameters and the shock normal direction remain consistent across each spacecraft crossing of the shock. The shock surface appears to be tilted almost the same degree as the Parker spiral, and the driver could be a Co–Rotating Interaction Region (CIR). During the April 23, 2007 event, the shock parameters do not change significantly except for the shock angle $\theta_{Bn}$, however, the shape of the IP shock appears to be twisted along the perpendicular direction to the Sun-Earth line as well. The driver of this rippled shock is Stream Interaction Region (SIR)/CIR as well.
@@ 1 Introduction
The solar corona is hotter than the photosphere, the chromosphere, and the transient layers beneath it. As a result, the high temperatures ionize atoms, creating a plasma of free-moving electrons and ions, the so-called solar wind. Historically, Parker (1958) predicted the existence of the solar wind and coined terms describing it. He deduced it based on German astronomer Ludwig Bierman’s observation of how the comet tail always points away from the Sun (Biermann, 1957). The existence of the solar wind was confirmed by the Mariner 2 spacecraft (Snyder & Neugebauer, 1965). The solar wind is a collisionless plasma, whose speed is both supersonic and super-Alfvenic. The latter means that its speed exceeds the Alfvén speed, which is a representative speed of magnetohydrodynamic waves in a plasma. Due to the abrupt change of the solar wind speed from supersonic to subsonic (e.g. by encountering an obstacle), a shock wave emerges. Interplanetary (IP) shock waves are common in the heliosphere, which is a bubble-like region of space surrounding the Sun and extending far beyond the orbits of the planets and is filled with the solar wind. There are many varieties of shocks such as planetary bow shocks, shocks that rise due to the stream interaction regions (SIR; see for example in Richardson, 2018; Koban et al., 2023; Silwal et al., 2024), and coronal mass ejection (CME) driven shocks. After completing a solar rotation, SIRs are referred to as Corotating Interaction Regions (CIRs). IP shocks are one of the main and efficient accelerators of energetic particles in the heliosphere (Tsurutani & Lin, 1985; Keith & Heikkila, 2021). These accelerated particles can enter the geomagnetic field and are hazardous to astronauts and satellites. IP shocks driven by CMEs can trigger large geomagnetic storms that can damage oil and gas pipelines and interfere with electrical power infrastructures (Gonzalez et al., 1994). GPS navigation and high-frequency radio communications are also affected by ionosphere irregularities brought on by geomagnetic storms (Cid et al., 2014) and can cause internet disruptions around the world (Jyothi, 2021; Facsko et al., 2024a). Therefore, IP shocks are important in determining and understanding space weather.
@@
Some dayside magnetospheric transient events (Zhang et al., 2022, see references therein) are generated by the interaction of a discontinuity frozen in the solar wind and the terrestrial bow shock. This fact raises the questions of how stable the discontinuities are, how long these phenomena travel with the solar wind, and whether they evolve during their travel, or show some wave dispersion (Facskó et al., 2005, István Ballai, personal communication). Dayside magnetospheric transient events are generated by tangential (TD) and rotational discontinuities (RD). However, there are catalogues for neither TDs nor RDs. Therefore, it is easier to study IP shocks because many were observed and listed previously (see Section 3).
@@
Reshetnyk et al. (2010) studied the 2D curvature of two IP shock events. Reshetnyk et al. (2010)’s paper showed that we do not know so much about the shape of discontinuities, especially about those of IP shocks. Facsko et al. (2010) extended the previous study to 3D. More events and more discontinuities (IP shocks) were needed. A special automated shock identification method was developed (Facsko et al., 2024b), however, finally, the appropriate events to study the temporal development and 3D shape of IP shocks were found in an open database developed by Nikbakhsh (2014); Kilpua et al. (2015). Finally, the origin of the IP shocks is also very interesting. Based on our current knowledge the shock could be CME driven, the result of SIR/CIR development, or caused by a flare. The determination of the 3D shape of IP shocks with a backward projection might explain the origin of these very important events.
@@
The shape of an IP shock front depends on the inhomogeneities of the solar wind through which the shock is propagating (Heinemann & Siscoe, 1974; Odstrčil & Pizzo, 1999; Dryer et al., 2004). Russell et al. (1983a, b) studied IP shock using two- and four-spacecraft measurements assuming that the IP shocks were planar locally. Terasawa et al. (2005) observed $\mathrm{15-20}^{\circ}$ shock surface ripple using Geotail (Nishida, 1994) and Advanced Composition Explorer (ACE; Stone et al., 1998) data. Szabo (2005) reported discrepancies of IP shock normal directions observing the same IP shock using ACE and Wind (Lin et al., 1995; Wilson et al., 2021), or Wind and IMP–8 (Team, 1974) spacecraft. Lepping et al. (2003) determined IP shock normals using Wind and IMP–8 observations at 1 AU (Astronomical Unit $\approx 1.5\times 10^{8}\,\mathrm{km}$). The average radius of the IP shock curvature was a few million km. Neugebauer & Giacalone (2005) collected and analyzed 26 IP shocks between the Earth’s bow shock and $L_{1}$ Lagrangian point. The average radius of the IP shock curvatures was a bit larger than Lepping et al. (2003)’s result, however, it also was a few million km. This result agreed with previous simulation results (Neugebauer & Giacalone, 2005).
@@
The Helios probe (Howard, 1975) observed IP shocks to evolve and change their behaviour at various heliospheric distances in the inner heliosphere (Lai et al., 2012) and between 1 and 5 AU by the Ulysses mission (Wenzel et al., 1992; Burton et al., 1992; Zhao et al., 2018). Trotta et al. (2023) studied shocklets using ACE, Wind, DSCOVR (Deep Space Climate Observatory), Time History of Events and Macroscale Interactions during Substorms (THEMIS, Angelopoulos, 2008) B and C measurements. Trotta et al. (2024) observed a CME driven shock at 0.07 AU and 0.7 AU using the Parker Solar Probe (PSP, Fox et al., 2016), and the Solar Orbiter (SolO, Marsch et al., 2005) observations. The PSP crossed a thick magnetosheath at the flank according to the early evolution phase of the CME. The SolO observed high energy (2 MeV) shock accelerated protons and a much more disturbed shock environment than the one observed by PSP. Therefore, the local small-scale environment of the shock was very different at different locations of the heliosphere.
In this study, we focus on the macroscopic structure and parameters of the observed IP shock events. The structure of this paper is as follows: we first list and describe briefly the missions, instruments, and databases that we used in this study in Section 2. We present the analyzed IP shock events in Section 3, then we describe the applied analysis methods in Section 4. Then we introduce the datasets that we analysed in Section 5. We discuss and analyse the observations in Section 6. Finally, we summarize the results of our study in Section 7.
@@ 2.1 The STEREO mission
The PLAsma and SupraThermal Ion Composition (PLASTIC) instrument measures protons and alpha particle parameters, as well as the composition of heavy ions in the solar wind plasma (Galvin et al., 2008). The In-situ Measurements of Particles and CME Transients magnetic field experiment (IMPACT) suite of instruments consists of seven instruments among which three are located on a 6-meter long boom. The other four are installed in the main part of the spacecraft (Acuña et al., 2008). The IMPACT measures the parameters of protons, heavy ions, and electrons, and the MAG magnetometer sensor in it measures the in situ magnetic fields in a range of ±$\pm 512\,nT$ with 0.1 nT resolution (Kaiser & Adams, 2007).
We downloaded all data from NASA’s Coordinated Data Analysis Web (CDAWeb, https://cdaweb.gsfc.nasa.gov/). We obtained the STEREO--A and B magnetic field observation and the ion plasma data from the STEREO IMPACT with time resolution of 100 ms and the STEREO PLASTIC instruments with time resolution of 1 minute, respectively. The magnetic field and the plasma data of the STEREO--A and B are in the RTN spacecraft coordinate system, where R is radially outward from the Sun, T is along the planetary orbital, and N is northward direction.
@@ 2.2 The Wind mission
NASA’s Wind spacecraft was launched on November 1, 1994, (Lin et al., 1995; Wilson et al., 2021). The Wind was initially sent to $L_{1}$ Lagrange point, however, the reaching of $L_{1}$ was delayed to be able to study in advance the terrestrial magnetosphere and lunar environment, as well. Following a sequence of orbital adjustments, the Wind spacecraft was positioned in a Lissajous orbit close to the $L_{1}$ Lagrange point in early 2004 for studying the incoming solar wind (NASA WIND team, 2020).
The spacecraft is equipped with eight instruments, however, we used only the Magnetic Field Investigation (MFI; Lepping et al., 1995) and the Three-Dimensional Plasma and Energetic Particle Investigation (3DP Lin et al., 1995) measurements in this study. The MFI consists of two magnetometers at the 12-meter boom, they measure magnetic field from 4 nT, to 65536 nT with a time resolution of 22 or 11 vectors per second in the calibrated high-resolution mode, while in the primary science mode the resolution can be three seconds, one minute and one hour (NASA WIND team, 2022). The 3DP measures the solar wind key parameters such as velocity, density, and temperature.
From CDAWeb, we got Wind magnetic and plasma data from the MFI and 3DP instruments with a time resolution of 3 s, for both. The magnetic field and the plasma data of the Wind are in the Geocentric Solar Ecliptic System (GSE) coordinate system, where the X-axis is pointing to the Sun from the Earth, Y-axis is in the ecliptic plane showing opposite to the planetary motion, and Z-axis is the northward direction.
@@ 2.3 The ACE mission
NASA’s ACE spacecraft was launched on August 25, 1997, (Stone et al., 1998). The spacecraft is located at $L_{1}$ Lagrangian point similarly to the Wind spacecraft. The spacecraft is equipped with nine primary scientific instruments and one engineering instrument. We used in the study the measurements of the Magnetometer (MAG; Smith et al., 1998), and the Solar Wind Electron, Proton and Alpha Monitor (SWEPAM; McComas et al., 1998). The MAG consists of twin triaxial flux-gate magnetometers such that magnetometer sensors have between 3 and 6 vectors $s^{-1}$ resolutions for continuous observation of the interplanetary magnetic field (Smith et al., 1998).
@@ 2.4 The Cluster fleet
The FGM is composed of two tri-axial fluxgate magnetometers, which are installed on one of the two 5-meter radial booms. It measures in the dynamic range of ±$\pm 65,536$ nT. At the highest dynamic level, the resolution is ±$\pm 8$ nT, and the time resolution is 100 vectors per second (Balogh et al., 1997, 2001; Organization, 2021). The CIS instrument measures three-dimensional ion distribution, and it is composed of two distinct sensors: the Composition Distribution Function (CODIF) sensor and the Hot Ion Analyzer (HIA) sensor (Rème et al., 1997, 2001). The CIS experiment is not operational for Cluster-2 and the HIA sensor is switched off for Cluster-4 due to a problem with the high voltage of the electrostatic analyzer (CIS team, 2021). Hence, for Cluster-2 and Cluster-4, the EFW instrument measurement became more indispensable for the analyses because we could obtain electron density in another way, independently from ion and electron plasma measurements (Gustafsson et al., 1997, 2001).
@@ 3 IP shock events and transformations
IP shock candidates are chosen from the shock lists in the Database of Heliospheric Shock Waves maintained at the University of Helsinki http://www.ipshocks.fi/. For the event selection, we chose the year 2007 because STEREO--A and STEREO--B were close to each other near the Sun-Earth line. In the first event, May 07 2007, the selected spacecraft are STEREO--A, STEREO--B, Wind, and the four Cluster satellites. For the second event, April 23, 2007, the spacecraft are STEREO--A and B, ACE, and Wind.
After choosing the shock candidates and obtaining the data, we transformed data coordinates in such a way that all the coordinate systems are changed to the Heliocentric Earth Ecliptic (HEE) coordinate system, where the X-axis is toward the Earth from the Sun, Z-axis is the ecliptic northward direction, while Y axis completes the right-handed Cartesian system. To transform from the RTN and the GSE coordinate systems to the HEE coordinate system, we used Transformation de REpères en Physique Spatiale (TREPS; http://treps.irap.omp.eu/) online tool. The TREPS tool, which is developed by the French Plasma Physics Data Centre (CDPP), the national data centre in France for the solar system plasmas, is based on SPICE (Spacecraft, Planet, Instrument, C-matrix, and Events) information system kernels created by National Aeronautics and Space Administration (NASA)/Navigation and Ancillary Information Facility (NAIF) tool (Génot et al., 2018).